OpenStudy (anonymous):

solve by substitution 7x+4y=10, 52-9y=x

OpenStudy (anonymous):

7(52-9y) + 4y = 10 364-63y + 4y = 10 354= 59y y= 6

OpenStudy (anonymous):

And x = 52 - 9y = 52-54 =-2